# Multi-view Graph Structural Representation Learning via Graph Coarsening

### v1

@@ 5.1 Experimental Settings
Table 1: The statistics of datasets from TUDataset. Avg.N and Avg.E denote the average number of nodes and edges.
Method CO IB IM DD N1 PTC
GK 72.84$\pm$0.28 65.87$\pm$0.98 43.89$\pm$0.38 78.45$\pm$0.26 62.30$\pm$0.30 57.26$\pm$1.41
WL 79.02$\pm$1.77 73.40$\pm$4.63 49.33$\pm$4.75 79.78$\pm$0.36 82.19$\pm$0.18 57.97$\pm$0.49
GAT 75.80$\pm$1.60 70.50$\pm$2.30 47.80$\pm$3.10 - 75.20$\pm$3.30 66.70$\pm$5.10
GMT 80.74$\pm$0.54 73.48$\pm$0.76 50.66$\pm$0.82 78.72$\pm$0.59 - -
MSGNN - 79.21$\pm$1.07 57.38$\pm$1.09 87.29$\pm$2.17 76.89$\pm$3.23 -
PSCN 72.60$\pm$2.15 71.00$\pm$2.29 45.23$\pm$2.84 77.12$\pm$2.41 76.34$\pm$1.68 62.29$\pm$5.68
GraphSAGE 79.70$\pm$1.70 72.40$\pm$3.60 49.90$\pm$5.00 65.80$\pm$4.90 - 63.90$\pm$7.70
DGCNN 73.76$\pm$0.49 70.03$\pm$0.86 47.83$\pm$0.85 79.37$\pm$0.94 74.44$\pm$0.47 58.59$\pm$2.47
CapsGNN 79.62$\pm$0.91 73.10$\pm$4.83 50.27$\pm$2.65 75.38$\pm$4.17 78.35$\pm$1.55 66.00$\pm$1.80
SEK-GIN - 76.90$\pm$3.80 53.80$\pm$3.20 - - 66.86$\pm$7.20
InfoGraph 70.65$\pm$1.13 73.03$\pm$0.87 49.69$\pm$0.53 72.85$\pm$1.78 73.80$\pm$0.70 61.65$\pm$1.43
GraphACL 74.72$\pm$0.55 74.53$\pm$0.39 51.65$\pm$0.34 79.34$\pm$0.42 - -
CGKS 76.80$\pm$0.10 - - - 79.10$\pm$0.20 63.50$\pm$1.30
U2GNN 77.84$\pm$1.48 77.04$\pm$3.45 53.60$\pm$3.53 80.23$\pm$1.48 69.81$\pm$2.36 69.63$\pm$3.60
MSLgo 97.60$\pm$0.81 88.40$\pm$2.73 63.60$\pm$2.83 93.21$\pm$3.26 79.34$\pm$1.88 71.53$\pm$6.59
@@
Table 2: The main results of the graph classification task on six datasets (mean accuracy (%) and standard deviation). The best scores are in bold. ’-’ indicates that results are unavailable in the original or published papers.
Baselines.
To evaluate the effectiveness of our model, we select 14 related works from 5 categories for comparison. (I) 2 kernel-based methods: GK Shervashidze et al. (2009) and WL Shervashidze et al. (2011); (II) 3 attention-based and pooling methods: GAT Velickovic et al. (2018), GMT Baek et al. (2021), and MSGNN Lv et al. (2023); (III) 5 GNN-based methods: PSCN Niepert et al. (2016), GraphSAGE Hamilton et al. (2017), DGCNN Zhang et al. (2018), CapsGNN Xinyi and Chen (2019), and SEK-GIN Yao et al. (2023); (IV) 3 contrastive learning methods: InfoGraph Sun et al. (2020), GraphACL Luo et al. (2023), and CGKS Zhang et al. (2023b); and (V) 1 backbone: U2GNN Nguyen et al. (2022) with Transformer style.
Implementation Details.
@@
We follow the work Nguyen et al. (2022) to evaluate the performance of our proposed method, adopting accuracy as the evaluation metric for graph classification tasks. To ensure a fair comparison between methods, we use the same data splits and 10-fold cross-validation accuracy to report the performance. In detail, we set the batch size to 4 and the dropout to 0.5. We pick the hierarchy depth in $\{1,2,3\}$, the number of Transformer Encoder layers in $\{1,2,3,4\}$, the number of sampling neighbors in $\{4,8,16\}$, the hidden size in $\{128,256,512,1024\}$, and the initial learning rate in $\{5e^{-4},1e^{-4},5e^{-4},1e^{-3}\}$. We utilize Adam Kingma and Ba (2015) as the optimizer. All experiments are trained and evaluated on an NVIDIA RTX 3050 OEM GPU.
(a) Boxplot and Loss
(b) 10 folds in last epoch
Figure 3: Robustness studies on the CO dataset.
5.2 Results and Analysis
Main Results.
@@ 5.2 Results and Analysis
Table 2 shows the main results of the graph classification task. MSLgo outperforms all other baseline models on all datasets except N1, where we achieve a suboptimal performance. In detail, the boost of MSLgo is higher in social networks than in bioinformatics. The main reason is that cliques of social networks are more distinct, while plain cliques and loop structures can not cover functional molecules and proteins completely. The improvement of PTC is only 1.90%, which concludes that long chain structures make the compression harder, where fewer loops and cliques appear.
Robustness Studies.
We also investigate the robustness of MSLgo. Figure 3 shows the boxplot of the performance for 10-fold and the average loss under each epoch of training on the CO dataset. It indicates that the performance stabilizes as the training progress. On the one hand, for every 10-fold, the standard deviation of the performance gradually decreases, especially for the last 5 epochs, where the accuracy stays above 95%. On the other hand, there are outliers (mainly lower performance) in some epochs. This phenomenon only occurs in the first half of training and disappears after Epoch 26, which explains that the learning ability of MSLgo becomes stable and no longer fluctuates.
Method CO IB IM
MSLgo 97.60$\pm$0.81 88.40$\pm$2.73 63.60$\pm$2.83
w/o LG 97.46$\pm$0.57 87.60$\pm$3.20 60.73$\pm$3.68
w/o GC 89.66$\pm$1.56 82.60$\pm$3.10 59.40$\pm$3.99
U2GNN 77.84$\pm$1.48 77.04$\pm$3.45 53.60$\pm$3.53
Method DD N1 PTC
MSLgo 93.21$\pm$3.26 79.34$\pm$1.88 71.53$\pm$6.59
w/o LG 91.43$\pm$2.77 78.73$\pm$1.28 71.25$\pm$5.39
w/o GC 83.78$\pm$4.63 72.38$\pm$2.14 70.67$\pm$5.21
U2GNN 80.23$\pm$1.48 69.81$\pm$2.36 69.63$\pm$3.60
Table 3: Ablation studies on all six datasets, social (upper) and bioinformatics (below).
5.3 Ablation Study
@@
5.4 Hyper-parameters Study

### v2

@@ 5.1. Experimental Settings
Table 1. The statistics of datasets from TUDataset. Avg.N, Avg.E, and Avg.D denote the average number of nodes, edges, and degrees.
Dataset # Graphs # Classes Avg.N Avg.E Avg.D Category
CO 5000 3 74.49 2457.78 37.39 Social Network
IB 1000 2 19.77 96.53 8.89 Social Network
IM 1500 3 13.00 65.94 8.10 Social Network
DD 1178 2 284.32 715.66 4.98 Bioinformatics
N1 4110 2 29.87 32.30 2.16 Molecules
PTC 344 2 25.56 25.96 1.99 Bioinformatics
PRO 1113 2 39.06 72.82 3.73 Bioinformatics
N109 4127 2 29.68 32.13 2.16 Molecules
We evaluate our approach on eight widely used datasets from TUDataset (Morris et al., 2020), including three social network datasets: COLLAB (CO), IMDB-BINARY (IB), and IMDB-MULTI (IM), two molecules datasets: NCI1 (N1) and NCI109 (N109), and three bioinformatics datasets: D&D (DD), PTC_MR (PTC), and PROTEINS (PRO). Table 1 shows the details of the datasets.
  • COLLAB is a scientific collaboration dataset in physics. A graph corresponds to a researcher’s ego network, which has three possible labels indicating the fields that the research belongs to.
  • IMDB-BINARY (MULTI) shows a movie collaboration network between actors/actresses and movies in IMDB. An edge connects two people when they participate in the same scene. The labels of BINARY are Action and Romance, and MULTI’s are Comedy, Romance, and Sci-Fi.
  • NCI1 (109) is relative to anti-cancer screens. Each input graph represents a chemical compound, treating atoms and bonds as nodes and edges. Classification tasks reveal whether the chemicals appear positive or not. NCI1 aims for non-small cells and NCI109 for ovarian cells.
  • D&D is a dataset of proteins based on the same data as PROTEINS, where nodes represent secondary structure elements with their type information and edges show spatial proximity.
  • PTC_MR reports the carcinogenicity for rats of 344 chemical compounds.
  • PROTEINS is a dataset of proteins. In each graph, nodes represent the amino acids, and edges exist if two nodes are less than 6 Angstroms apart. It includes two categories: enzymes and non-enzymes.
@@
Table 2. The main results of the graph classification task on eight datasets (mean accuracy (%) and standard deviation). The best scores are in bold and underline. ’-’ indicates that results are unavailable in the original or published papers. A.R shows the average ranking of each method. We highlight every method with outstanding performance in each category.
Method CO IB IM DD N1 PTC PRO N109 A.R
GW $72.84^{0.28}$ $65.87^{0.98}$ $43.89^{0.38}$ $78.45^{0.26}$ $62.30^{0.30}$ $57.26^{1.41}$ $71.67^{0.55}$ 18.71
WL $79.02^{1.77}$ $73.40^{4.63}$ $49.33^{4.75}$ $79.78^{0.36}$ $\underline{\bf{82.19}}^{0.18}$ $57.97^{0.49}$ $74.68^{0.49}$ $\underline{\bf{82.46}}^{0.24}$ 8.75
GCN $81.72^{1.64}$ $73.30^{5.29}$ $51.20^{5.13}$ $79.12^{3.07}$ $76.00^{0.90}$ $59.40^{10.30}$ $75.65^{3.24}$ $67.09^{3.43}$ 9.13
GAT $75.80^{1.60}$ $70.50^{2.30}$ $47.80^{3.10}$ $74.40^{0.30}$ $74.90^{0.10}$ $66.70^{5.10}$ $74.70^{2.20}$ 14.57
GraphSAGE $79.70^{1.70}$ $72.40^{3.60}$ $49.90^{5.00}$ $65.80^{4.90}$ $63.90^{7.70}$ $65.90^{2.70}$ $64.67^{2.41}$ 14.67
DGCNN $73.76^{0.49}$ $70.03^{0.86}$ $47.83^{0.85}$ $79.37^{0.94}$ $74.44^{0.47}$ $58.59^{2.47}$ $75.54^{0.94}$ 14.00
CapsGNN $79.62^{0.91}$ $73.10^{4.83}$ $50.27^{2.65}$ $75.38^{4.17}$ $78.35^{1.55}$ $66.00^{1.80}$ $76.28^{3.63}$ 9.57
GIN $80.20^{1.90}$ $75.10^{5.10}$ $52.30^{2.80}$ $75.20^{2.90}$ $79.10^{1.40}$ $64.60^{7.00}$ $76.20^{2.80}$ $68.44^{1.89}$ 7.75
GDGIN - - - $77.80^{3.60}$ $60.30^{4.50}$ $73.70^{3.40}$ 15.33
GraphMAE $80.32^{0.46}$ $75.52^{0.66}$ $51.63^{0.52}$ $80.40^{0.30}$ $75.30^{0.39}$ 6.00
InfoGraph $70.65^{1.13}$ $73.03^{0.87}$ $49.69^{0.53}$ $72.85^{1.78}$ $73.80^{0.70}$ $61.65^{1.43}$ $74.40^{0.30}$ 15.14
JOAO $75.53^{0.18}$ $70.21^{3.08}$ $47.22^{0.41}$ $75.81^{0.73}$ $74.86^{0.39}$ $73.31^{0.48}$ 17.50
RGCL $70.92^{0.65}$ $71.85^{0.84}$ $78.86^{0.48}$ $75.03^{0.43}$ 15.75
CGKS $76.80^{0.10}$ $79.10^{0.20}$ $63.50^{1.30}$ $76.00^{0.20}$ 7.75
DIFFPOOL $75.48^{0.00}$ $73.14^{0.70}$ $49.53^{3.98}$ $80.64^{0.00}$ $62.32^{1.90}$ $76.25^{0.00}$ $61.98^{1.98}$ 11.00
SAGPool $78.85^{0.56}$ $72.55^{1.28}$ $49.33^{4.90}$ $76.45^{0.97}$ $74.18^{1.20}$ $71.86^{0.97}$ $74.06^{0.78}$ 14.71
ASAP $78.64^{0.50}$ $72.81^{0.50}$ $50.78^{0.75}$ $76.87^{0.70}$ $71.48^{0.42}$ $74.19^{0.79}$ $70.07^{0.55}$ 13.43
GMT $80.74^{0.54}$ $73.48^{0.76}$ $50.66^{0.82}$ $78.72^{0.59}$ $75.09^{0.59}$ 9.60
SAEPool - - - - $74.48^{0.00}$ $80.36^{0.00}$ $75.85^{0.00}$ 6.67
MVPool - $52.00^{0.80}$ $82.20^{1.40}$ $80.10^{1.30}$ $\underline{\bf{85.70}}^{1.20}$ $81.90^{1.60}$ 6.20
PAS - - $53.13^{4.49}$ $79.62^{1.75}$ $77.36^{3.69}$ 5.33
GraphGPS - - - - $79.44^{0.65}$ $53.75^{6.20}$ $76.27^{0.95}$ 12.67
U2GNN $77.84^{1.48}$ $77.04^{3.45}$ $53.60^{3.53}$ $80.23^{1.48}$ $69.63^{3.60}$ $78.53^{4.07}$ 4.17
UGT - - - - $77.55^{0.16}$ $80.12^{0.32}$ $75.45^{1.26}$ 6.33
CGIPool $80.30^{0.69}$ $72.40^{0.87}$ $51.45^{0.65}$ $78.62^{1.04}$ $74.10^{2.31}$ $77.94^{1.37}$ 9.83
DosCond - - - $78.92^{0.64}$ $71.70^{0.20}$ 13.00
KGC - $69.20^{1.37}$ $61.58^{2.49}$ $66.43^{0.92}$ 18.33
SS - $73.89^{0.70}$ $51.80^{1.20}$ $79.78^{0.49}$ $76.13^{0.60}$ 6.00
GRLsc $\underline{\bf{98.18}}^{0.42}$ $\underline{\bf{89.40}}^{2.50}$ $\underline{\bf{63.20}}^{3.55}$ $\underline{\bf{99.58}}^{0.78}$ $79.08^{1.92}$ $\underline{\bf{76.43}}^{5.22}$ $76.37^{4.37}$ $76.98^{1.90}$ 2.63
5.1.2. Baselines.
To fully evaluate the effectiveness of our model, we select 28 related works from 6 categories for comparison.
As for the Graph Transformer framework, we pick 17 baselines. They are (I) 2 kernel-based methods: GK (Shervashidze et al., 2009) and WL (Shervashidze et al., 2011); (II) 8 GNN-based methods: GCN (Kipf and Welling, 2017), GAT (Velickovic et al., 2018), GraphSAGE (Hamilton et al., 2017), DGCNN (Zhang et al., 2018), CapsGNN (Xinyi and Chen, 2019), GIN (Xu et al., 2019), GDGIN (Kong et al., 2022), and GraphMAE (Hou et al., 2022); (III) 4 Contrastive Learning methods: InfoGraph (Sun et al., 2020), JOAO (You et al., 2021), RGCL (Shuai et al., 2022), and CGKS (Zhang et al., 2023b); and (IV) 3 GT-based methods: GraphGPS (Rampásek et al., 2022), U2GNN (Nguyen et al., 2022), and UGT(Hoang and Lee, 2024).
We also choose 11 models from 2 categories to verify our Graph Coarsening technique. One is (V) 7 Graph Pooling methods: DIFFPOOL (Ying et al., 2018), SAGPool(Lee et al., 2019), ASAP (Ranjan et al., 2020), GMT (Baek et al., 2021), SAEPool (Zhu et al., 2022), MVPool (Zhang et al., 2023a), and PAS (Wei et al., 2024). The other is (VI) 4 Graph Coarsening and Condensing methods: CGIPool (Pang et al., 2021), DosCond (Jin et al., 2022a), KGC (Chen et al., 2023c), and SS (Zhang and Zhao, 2024).
To further explain the unique design of our LCC block, we select 3 coarsening techniques: networkx (Hagberg et al., 2008), KGC (Chen et al., 2023c), and L19 (Loukas, 2019). Each one takes two variants: neighborhood and clique. We will explain the details later in Section 5.4.
5.1.3. Implementation Details.
@@
We follow the work (Xu et al., 2019; Nguyen et al., 2022) to evaluate the performance of our proposed method, adopting accuracy as the evaluation metric for graph classification tasks. To ensure a fair comparison between methods, we use the same data splits and 10-fold cross-validation accuracy to report the performance. In detail, we set the batch size to 4 and the dropout to 0.5. We pick the hierarchy depth in $\{1,2,3\}$, the number of Transformer Encoder layers in $\{1,2,3,4\}$, the number of sampling neighbors in $\{4,8,16\}$, the hidden size in $\{128,256,512,1024\}$, and the initial learning rate in $\{5e^{-4},1e^{-4},5e^{-4},1e^{-3}\}$. We utilize Adam (Kingma and Ba, 2015) as the optimizer. All experiments are trained and evaluated on an NVIDIA RTX 3050 OEM 8GB GPU and 16GB CPU. Our code is available: https://github.com/NickSkyyy/LCC-for-GC.
5.2. Results and Analysis
5.2.1. Main Results.
@@ 5.2. Results and Analysis
Table 2 shows the main results of the graph classification task. GRLsc outperforms other baseline models on most datasets, with the highest average ranking 2.63 above all. In detail, the boost of GRLsc is higher in social networks than in molecules and bioinformatics. WL, GraphMAE, CGKS, PAS, U2GNN, and SS show competitive results among their categories, even better than GRLsc (e.g., WL in N1 and N109). The main reason is that cliques of social networks are more distinct, while plain cliques and loop structures can not cover functional molecules and proteins completely. These main experimental results show the limitation of our model: the classification ability slightly drops when fewer loops and cliques appear. N1 and N109 contain many simple loops and long chains, while some unconnected graphs in PRO make it hard to coarsen and convert. We will show more details later in the ablation of the LCC block (Section 5.4) and case study (Section 5.6).
5.2.2. Robustness Studies.
We also investigate the robustness of GRLsc. Figure  5 presents the model performance of 10-fold in the best epoch for each dataset. Most experiment folds are higher than or equal to the average, having a mean turbulence of 2.58%. The worst two are PTC and PRO, with standard deviations of 5.22% and 4.37%, respectively. In PTC, the performance gap between the best fold (1 and 4, 82.86%) and the worst fold (5 and 9, 67.65%) is 15.21%, about 5 in 34 graphs. The best two are CO and DD, with standard deviations of 0.42% and 0.78%, respectively. In CO, the performance gap between the best fold (1, 98.80%) and the worst fold (8, 97.20%) is 1.60%, about 8 in 500 graphs.
5.3. Block Ablation Study
In this section, we will evaluate the effectiveness of each component of our model. We design two sets of ablation experiments, one for modules to analyze the effectiveness of GRLsc. And the other to discuss the impact of different coarsening strategies on downstream classification tasks.
@@ 5.3. Block Ablation Study
Figure 5. 10 folds / best epoch on every dataset. The red dashed line plots the average accuracy.
Table 3. Block ablation studies on all eight datasets.
Method CO IB IM DD N1 PTC PRO N109
GRLsc $\underline{\bf{98.18}}^{0.42}$ $\underline{\bf{89.40}}^{2.50}$ $\underline{\bf{63.20}}^{3.55}$ $\underline{\bf{99.58}}^{0.78}$ $\underline{\bf{79.08}}^{1.92}$ $\underline{\bf{76.43}}^{5.22}$ $\underline{\bf{76.37}}^{4.37}$ $\underline{\bf{76.98}}^{1.90}$
- ${}_{w/o\,LGC}$ $97.96^{0.56}$ $87.60^{2.58}$ $61.47^{3.10}$ $99.32^{1.06}$ $77.42^{1.64}$ $69.50^{6.24}$ $75.02^{4.12}$ $74.90^{2.05}$
- ${}_{w/o\,LCC}$ $97.60^{0.81}$ $87.10^{2.91}$ $58.60^{3.74}$ $83.78^{4.63}$ $75.33^{2.12}$ $66.92^{8.38}$ $74.48^{3.42}$ $72.79^{2.14}$
- ${}_{w/o\,LGC+LCC}$ $96.50^{0.84}$ $84.10^{2.91}$ $57.87^{3.98}$ $83.02^{5.33}$ $72.53^{2.28}$ $63.09^{7.90}$ $68.01^{3.13}$ $69.20^{1.04}$
5.4. Coarsening Strategy Ablation
To better discuss the impact of different coarsening strategies on the model performance, we select four categories of graph coarsening schemes: random, networkx (Hagberg et al., 2008), KGC (Chen et al., 2023c), and L19 (Loukas, 2019). Under each category, there are two variants: neighbor and clique.
  • neighbor is an algorithm based on adjacent structures, which mainly aggregates information according to 1-hop adjacent nodes. There is no such algorithm in networkx, so we pick nx.cycle instead. It is similar to clique but based on cycle searches.
  • clique is an algorithm based on various clique searches, which aggregate node information according to clique partitions.
Table 4. Coarsening strategy ablation studies on all eight datasets.
Method CO IB IM DD N1 PTC PRO N109
Neighbor $98.12^{0.65}$ $89.10^{4.09}$ $62.80^{3.45}$ $98.64^{1.94}$ $77.79^{1.48}$ $70.67^{6.53}$ $75.20^{3.99}$ $76.40^{1.99}$
Random $97.94^{0.58}$ $89.10^{3.78}$ $61.70^{1.90}$ $99.32^{0.74}$ $77.79^{2.36}$ $71.81^{4.61}$ $72.51^{5.68}$ $74.10^{1.80}$
nx.cycle $98.02^{0.60}$ $87.00^{2.31}$ $62.47^{4.06}$ $86.16^{4.10}$ $74.84^{2.59}$ $75.92^{6.31}$ $74.30^{4.57}$ $73.81^{2.21}$
nx.clique $98.04^{0.58}$ $87.90^{2.62}$ $62.53^{3.41}$ $86.19^{5.62}$ $75.69^{1.92}$ $72.05^{6.67}$ $72.77^{6.16}$ $73.88^{1.81}$
KGC.neighbor $98.08^{0.58}$ $87.00^{1.95}$ $58.87^{4.03}$ $97.71^{2.57}$ $76.18^{2.12}$ $75.00^{2.66}$ $73.13^{3.70}$ $73.42^{1.36}$
KGC.clique $96.06^{0.77}$ $89.30^{2.72}$ $59.40^{3.66}$ $82.00^{7.04}$ $74.60^{1.91}$ $70.61^{8.86}$ $74.75^{5.05}$ $72.84^{1.72}$
L19.neighbor $96.92^{0.84}$ $87.10^{2.62}$ $61.70^{3.17}$ $96.44^{4.82}$ $75.94^{2.20}$ $74.06^{8.07}$ $74.21^{4.45}$ $73.81^{2.14}$
L19.clique $97.28^{0.66}$ $88.20^{2.79}$ $61.87^{3.55}$ $91.08^{3.00}$ $77.81^{1.48}$ $71.20^{6.10}$ $74.30^{3.65}$ $75.14^{1.86}$
LCC $\underline{\bf{98.18}}^{0.42}$ $\underline{\bf{89.40}}^{2.50}$ $\underline{\bf{63.20}}^{3.55}$ $\underline{\bf{99.58}}^{0.78}$ $\underline{\textbf{79.08}}^{1.92}$ $\underline{\bf{76.43}}^{5.22}$ $\underline{\textbf{76.37}}^{4.37}$ $\underline{\textbf{76.98}}^{1.90}$
In terms of implementation, we set $K=5$ and $r=0.5$, according to the paper (Chen et al., 2023c; Loukas, 2019). The algorithm strategy is greedy, and other parameters are consistent with the origin. To ensure alignment with GRLsc, the maximum coarsening level is set to 1, which means only coarsening the original graph without higher-level compressions. The random algorithm sets $R=K=5$, and each supernode randomly selects components from the sampling sets until all nodes are covered. Table 4 shows the model performance under different coarsening strategies, where LCC achieves the optimal results under all eight datasets. It indicates that the graph coarsening view obtained by LCC is more suitable for graph classification tasks.
5.4.1. Time Analysis
Figure 6. Runtimes of all coarsening algorithms. We make logarithmic processing due to different scales of datasets.
Table 5. Scale analysis of graph coarsening. Each group represents the size of the intermediate coarsening graph under different strategies, where $\bar{V}$ and $\bar{E}$ represent the average nodes and edges, and $r_{V}$ and $r_{E}$ represent the scaling ratio of the coarsening graph compared to the original graph. The negative number means reducing the scale from the original graph, and the positive means expanding.
Method CO IB IM DD N1 PTC PRO N109
Original $\bar{V}$ 74.49 19.77 13.00 284.32 39.06 25.56 29.87 29.68
$\bar{E}$ 2457.50 96.53 65.94 715.66 72.82 25.96 32.30 32.13
nx.cycle $\bar{V}$ 74.49 19.77 13.00 284.32 29.84 25.30 39.06 29.65
$r_{V}$ 0.00 0.00 0.00 0.00 -0.24 -0.01 +0.31 0.00
$\bar{E}$ 2474.63 101.15 67.94 722.44 42.09 47.85 77.02 41.98
$r_{E}$ +0.01 +0.05 +0.03 +0.01 -0.42 +0.84 +1.38 +0.31
nx.clique $\bar{V}$ 42.46 3.77 2.05 299.08 32.31 25.86 40.79 32.16
$r_{V}$0.43 -0.81 -0.84 +0.05 -0.17 +0.01 +0.37 +0.08
$\bar{E}$ 7082.97 11.96 2.81 2788.05 109.06 102.75 259.48 108.95
$r_{E}$ +1.88 -0.88 -0.96 +2.90 +0.50 +2.96 +7.03 +2.39
KGC.neighbor $\bar{V}$ 5.24 2.05 1.51 90.15 13.86 12.49 13.31 13.78
$r_{V}$0.93 -0.90 -0.88 -0.68 -0.65 -0.51 -0.55 -0.54
$\bar{E}$ 19.17 1.47 0.64 448.42 24.53 18.66 37.13 24.37
$r_{E}$0.99 -0.98 -0.99 -0.37 -0.66 -0.28 +0.15 -0.24
KGC.clique $\bar{V}$ 10.91 3.26 1.92 116.00 17.33 16.31 17.74 17.27
$r_{V}$0.85 -0.84 -0.85 -0.59 -0.56 -0.36 -0.41 -0.42
$\bar{E}$ 60.55 5.38 1.80 506.74 27.54 22.76 46.28 27.41
$r_{E}$0.98 -0.94 -0.97 -0.29 -0.62 -0.12 +0.43 -0.15
L19.neighbor $\bar{V}$ 46.30 12.46 9.83 142.45 15.27 13.67 19.88 15.17
$r_{V}$0.38 -0.37 -0.24 -0.50 -0.61 -0.47 -0.33 -0.49
$\bar{E}$ 1275.87 57.22 46.81 516.42 25.57 21.26 47.52 25.40
$r_{E}$0.48 -0.41 -0.29 -0.28 -0.65 -0.18 +0.47 -0.21
L19.clique $\bar{V}$ 46.12 12.42 9.76 142.45 17.35 16.00 20.25 17.28
$r_{V}$0.38 -0.37 -0.25 -0.50 -0.56 -0.37 -0.32 -0.42
$\bar{E}$ 1290.52 57.92 45.58 535.58 27.55 22.83 50.80 27.43
$r_{E}$0.47 -0.40 -0.31 -0.25 -0.62 -0.12 +0.57 -0.15
LCC $\bar{V}$ 12.52 3.45 2.01 155.72 18.88 20.49 25.85 18.73
$r_{V}$0.83 -0.83 -0.85 -0.45 -0.52 -0.20 -0.13 -0.37
$\bar{E}$ 29.38 3.07 1.22 306.33 18.97 20.07 40.54 18.85
$r_{E}$0.99 -0.97 -0.98 -0.57 -0.74 -0.23 +0.26 -0.41
We analyze the linear time complexity of LCC in section 4.2.4. Here, we compare it with other methods. Figure 6 shows the results.
The results show that the runtime of LCC is significantly lower than other methods. Except for CO and DD, the computation time is within 2s for all other datasets. Runtime for CO is 7.60s, while for DD is 37.58s. The average runtime of KGC.neighbor is 44.55s, and L19.clique is 54.98s. Both of them are higher than LCC.
5.4.2. Scale Analysis
Though we do not require LCC to achieve the optimal coarsening effect, we still analyze and compare the scale of the intermediate coarsening graph $G_{lcc}$. We experiment with three other categories of strategies and their variants in addition to random algorithms. Given a set of original input graphs $\mathcal{G}_{o}=\{G_{o}^{1},\cdots,G_{o}^{M}\}$, we calculate and collate the scale of $\mathcal{G}_{lcc}=\{G_{lcc}^{1},\cdots,G_{lcc}^{M}\}$. We show average node and edge numbers through $\bar{V}=\frac{1}{M}\sum_{i}\mathcal{V}_{G_{i}}$ and $\bar{E}=\frac{1}{M}\sum_{i}\mathcal{E}_{G_{i}}$, and analyze the scale changes leveraging formula $r_{V}=\frac{\bar{V}_{lcc}-\bar{V}_{o}}{\bar{V}_{o}}$ and $r_{E}=\frac{\bar{E}_{lcc}-\bar{E}_{o}}{\bar{E}_{o}}$. Table 5 shows the results.
(a) Loops: N109 (77)
(b) Cliques: IB (14)
(c) Plain: PTC (5)
Figure 7. Visualization of different coarsening strategies. For (a) loops, (b) cliques, and (c) plain graphs without loops and cliques, we find corresponding sample graphs from datasets (graph ID in parentheses). In each group, row 1 col 1 is the original input graph, row 2 col 1 is the coarsening result of GRLsc, and the other columns are two variants of the method in the same category. Different node colors represent different node classes. The graph layout does not contain semantics such as rotation and symmetry but only shows the connections.
In general, the LCC we designed can achieve a considerable coarsening result. The coarsening rate at the node level is about 52.2%. It is slightly higher at the edge level, 57.9% or so. In baselines, the best one is KGC.neighbor, with a 70.5% compression rate at the node level and 54.7% at the edge level. We do not expect LCC to achieve similar high rates. When the coarsening process is to the maximum extent, amounts of high-level structural information are lost.
Table 6. Optimal parameter settings.
Hyper-parameters CO IB IM DD N1 PTC PRO N109
$K$ 1 1 3 2 1 3 2 1
$T$ 2 2 1 2 2 4 1 1
$N$ 16 16 4 16 16 16 4 8
$|H|$ 1024 1024 1024 1024 1024 1024 1024 1024
$Lr$ 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
$|B|$ 4 4 4 4 4 4 4 4
Figure 8. Hyper-parameters Study. We evaluate four hyper-parameters: hierarchy depth $K$, the number of Transformer Encoder layers $T$, the number of sampling neighbors $N$, and the hidden size $|H|$.
As the number of nodes decreases linearly, edges have a faster decay. We cannot mine enough structural information in very simple hypergraphs. Thus, LCC achieves the maximum balance between coarsening nodes and preserving structures.
5.4.3. Visualization
We further analyze the coarsening comparison using visualization. Figure 7 shows the results.
Figures 7(a) and  7(b) visualize the loops and cliques that GRLsc focuses on. It appears that GRLsc can obtain the coarsening graph representation closest to the original graph. We retain the connections between supernodes to the greatest extent and mine the high-level structural information thoroughly based on the coarsening procedures. Other algorithms lack balance in coarsening and structural representation. Random algorithms break away from the original structural semantics and pursue to cover the whole graph by specifying the number of nodes. KGC presents a more concise coarsening pattern. However, it fails to express high-level structural information and has poor adaptability to loops. L19 relaxes the restriction, not limiting nodes located in multiple partitioning sets, but still cannot refine the structures.
In addition, we analyze for plain graphs that do not contain loops and cliques, as shown in Figure 7(c). GRLsc returns the same graph $G_{lcc}$ as the original input $G_{o}$ through LCC and makes a further conversion via LGC. In this kind of graph, KGC can effectively identify other structures as claw-like ones and realize their coarsening process.
5.5. Hyper-parameters Study